# Quantum sampling on a quantum annealer for large volumes in the strong coupling limit for gauge group U(3)

### v2

@@ III.4 Branching Strategy
Another promising method to correct the histograms that D-Wave provides to obtain valid results is to repeatedly branch into sub-branches, and statistically evaluate each branch to obtain expectation values for the chiral condensate and the energy. The essential idea of the branching strategy is to maximize quantum parallelization. While in the Metropolis-Hastings the same accept-reject step is required, the branching strategy performs accept-reject at each branch.
@@ V Conclusion
We developed an algorithm to simulate ${\rm U}(3)$ gauge theory in the strong coupling limit for large 2-dimensional and 4-dimensional volumes on a quantum annealer. In particular, we made use of an hybrid quantum/classical approach, whereby we determined the histograms in the dual variables relevant to the QUBO formalism via the D-Wave quantum annealer, and than ran the Metropolis-Hastings algorithm based on these histograms on a classical computer. We optimized the hyper parameters to reduce the compute time on the annealer, and found optimal values for annealing_time and chain_strength. Also, the penalty factor $p$ that balances the weight matrix and the constraint in the QUBO matrix, for which the quality of the histograms is best was determined to be $p=1$, resulting in large acceptance rates.
@@
Extending to ${\rm SU}(3)$ and including gauge corrections are another important next step towards a more realistic effective theory of lattice QCD. The first aspect includes baryons while the second aspect includes gluon propagation. The dual representation is well established for both extensions, and again can be mapped on binary vectors. However, these formulations will require more logical qubits. Still, it remains feasible to map the logical qubits required for the QUBO matrix to the physical qubits on the annealer. However, while the histograms required for this work, given square sampling, had $4^{4}$ different boundaries to classify the histograms, these generalizations will require up to $16^{4}$ different boundaries, for which the histograms can no longer be pre-computed on D-wave. Instead they need to be computed during Metroplis-Hastings in between the updates, whenever required. Studying more realistic effective theories will introduce two more physical parameters [14]: the quark chemical potential $\mu_{q}$, and the inverse gauge coupling $\beta$ that is related to the lattice spacing $a$. We will report on our findings using the same approach via histograms determined on the annealer in a subsequent publication.

### v3

@@ III.4 Branching Strategy
In the previous section we studied in detail how to obtain the correct equilibrium distribution from the histograms that D-Wave provides via the Metropolis-Hastings algorithm. Note that if the histograms were already perfect Boltzmann distribution, we could instead use a heatbath algorithm. Another promising method to obtain valid results that we briefly outline is to repeatedly branch into sub-branches, and statistically evaluate each branch to obtain expectation values for thermal observables, such as the chiral condensate and the energy. The essential idea of the branching strategy is to maximize quantum parallelization. While in the Metropolis-Hastings the same accept-reject step is required, the branching strategy performs accept-reject at each branch.
@@ V Conclusion
We developed an algorithm to simulate ${\rm U}(3)$ gauge theory in the strong coupling limit for large 2-dimensional and 4-dimensional volumes on a quantum annealer. In particular, we made use of an hybrid quantum/classical approach, whereby we determined the histograms in the dual variables relevant to the QUBO formalism via the D-Wave quantum annealer, and than ran the Metropolis-Hastings algorithm based on these histograms on a classical computer. We described the hyper parameters we used to reduce the compute time on the annealer, detailing how we numerically determined optimal values for annealing_time and chain_strength by maximizing the validity rate given our finite compute time resources. Also, the penalty factor $p$ that balances the weight matrix and the constraint in the QUBO matrix, for which the quality of the histograms is best was determined to be $p=1$, resulting in large acceptance rates.
@@
Extending to ${\rm SU}(3)$ and including gauge corrections are another important next step towards a more realistic effective theory of lattice QCD. The first aspect includes baryons while the second aspect includes gluon propagation. The dual representation is well established for both extensions, and again can be mapped on binary vectors. However, these formulations will require more logical qubits. Still, it remains feasible to map the logical qubits required for the QUBO matrix to the physical qubits on the annealer. However, while the histograms required for this work, given square sampling, had $4^{4}$ different boundaries to classify the histograms, these generalizations will require up to $16^{4}$ different boundaries, for which the histograms can no longer be pre-computed on D-Wave. Instead they need to be computed during Metropolis-Hastings in between the updates, whenever required. Studying more realistic effective theories will introduce two more physical parameters [20]: the quark chemical potential $\mu_{q}$, and the inverse gauge coupling $\beta$ that is related to the lattice spacing $a$. We will report on our findings using the same approach via histograms determined on the annealer in a subsequent publication.